# Contract Design for Pandora’s Box

### v1

@@ 1 Introduction
In many real-world situations, e.g., the search for a job candidate or house, a decision maker is faced with the choice between different alternatives of a-priori unknown value, which can be explored at some cost. A large class of such settings can be modeled by the famous Pandora’s box model proposed and analyzed by Weitzman [36]. In this model, there are $n$ alternatives or boxes. The value in each box is distributed independently according to a known probability distribution. There is a known additive function for the exploration cost of the boxes. The goal is to maximize the value of the selected box minus the cost of explored boxes. Starting from [31], this problem has recently received a lot of attention at the intersection of economics and computer science (see, e.g., [35, 24, 7, 9, 6, 23])
@@
Due to resource and time constraints, in many markets the aforementioned exploration tasks are not executed directly by the decision maker. Rather, a decision maker (called principal $\mathcal{P}$ in the following) can delegate the exploration to an agent $\mathcal{A}$ (say, a consultant, head hunter, real estate agent, etc.), whose incentives are potentially misaligned with that of $\mathcal{P}$. As a consequence of delegating the search, $\mathcal{P}$ can merely observe the outcome of the search, not the other actions taken by $\mathcal{A}$. This leads to a possibly undesirable outcome for $\mathcal{P}$, even when they have the power to commit to accepting or rejecting certain outcomes [5].
@@
A natural approach to align the incentives of an agent with the goals of the principal in such hidden-action settings are contracts. A contract allows for utility transfers to be made from the principal to the agent as a function of the outcome of the delegated task. While contract theory is a well-established [8] and celebrated [32] area in economics, the interest in algorithmic contract design has surged only relatively recently, see e.g., [19], [20], [16], [13], [25].
@@
In this work, we initiate the study of algorithmic contract design for the Pandora’s Box problem. Specifically, we consider the following problem (for a fully formal definition, see Section 2): There are $n$ boxes, the $i$-th of which contains a prize, drawn independently from a known probability distribution, and can be opened at known cost $c_{i}$. The $j$-th prize from box $i$ has some utilities $a_{ij}$ and $b_{ij}$ for $\mathcal{A}$ and $\mathcal{P}$, respectively. Based on that information, $\mathcal{P}$ commits to a contract, specifying a transfer $t_{ij}\in[0,b_{ij}]$ for each prize, where we assume non-negative transfers (limited liability). Given such a contract, $\mathcal{A}$ faces an instance of the classic Pandora’s Box Problem, which they solve optimally. Assuming $\mathcal{A}$ selects prize $j$ from box $i$, the resulting final utility of $\mathcal{P}$ is $b_{ij}-t_{ij}$. The goal is to efficiently compute a contract that maximizes the expectation of the latter quantity. Notably, ties in the optimization of $\mathcal{A}$ are broken in favor of $\mathcal{P}$, which is a standard assumption that makes the maximization problem well-defined.
@@ 1.1 Our Contribution
With a solution of $\mathcal{A}$’s problem at hand, the problem of finding an optimal contract can be understood as follows: On the one hand, transfers manipulate the fair caps, the resulting exploration order, and the acceptance decisions of $\mathcal{A}$ in such a way that leads to a more favorable outcome for $\mathcal{P}$. On the other hand, transfers make $\mathcal{P}$ lose utility since they are costly. That trade-off needs to be solved.
We start by considering linear contracts [28, 19], a class of contracts that has been particularly successful due to its simplicity. Here, a contract is fully described by a number $\alpha\in[0,1]$, which defines transfers $t_{ij}=\alpha\cdot b_{ij}$. We identify a polynomial-time computable set of critical values of $\alpha$ at which the set of $\mathcal{A}$-optimal policies (i.e., the set of policies that satisfy the above description) changes. By reasoning about the behavior of $\mathcal{P}$’s utility between any two such values, we can argue that the optimal contract needs to be a critical value. This type of argument is reminiscent of arguments from the recent literature (e.g., [16], [21], [15]), but here it requires special care. In particular, due to the tie breaking in favor of $\mathcal{P}$, which—as discussed above—is very relevant here, we need to keep track of the set of $\mathcal{A}$-optimal policies.
@@
The final case we consider has a more general value structure, but we compromise on the boxes having different distributions. Specifically, we consider instances with only single positive prize for $\mathcal{P}$, and all distributions are identical. Still, transfers can be made depending on the box from which the prize originates. As for binary boxes, the complexity is reduced by the fact that only a single payment per box has to be decided. Through an intricate sequence of exchange arguments and the help of continuization, we establish that there is an optimal contract with a relatively simple structure: (i) In a first phase, the prize with positive principal value gets immediately accepted, and all transfers are identical; (ii) all transfers for boxes in the second phase are also identical. These conditions allow to find the optimal contract by enumerating a polynomial-sized set of contracts.
@@
Overview
After a review of related work in the subsequent section, we formally introduce our problem in Section 2 along with preliminary observations. Section 3 then treats linear contracts, and Section 4 deals with general contracts in the different special cases (no agent value in Section 4.1, binary boxes in Section 4.2, and i.i.d. with single positive prize in Section 4.3).
@@ 1.2 Related Work
Contract design has recently gained a lot of attention from a computational point of view. Duetting et al. [19] advocated the study of linear contracts as an alternative to the more complicated (and sometimes unintuitive) general contracts. They show robustness results and give parameterized approximation guarantees w.r.t. optimal (general) contracts. For settings in which the outcome space is succinctly represented, hardness results are shown in [20]. In a similar but different approach, combinatorial contracts [16] let the agent choose a subset of actions which stochastically determines the (binary) outcome. Important recent work in the area of combinatorial contracts includes [21], [15]. Contracts were also studied with multiple agents [17, 15], ambiguity [18], and private types [2] with connections to classic mechanism design [1]. More generally, contracts have been of interest in specific application domains, such as classification tasks in machine learning tasks [33].
@@
Kleinberg and Kleinberg [30] also study a variant with exploration costs by the agent. Their scenario assumes binary outcomes and is very similar to our setting with binary boxes, since the agent also faces an instance of the Pandora’s box problem. The key difference is, as for other delegation models, that $\mathcal{P}$ does not use transfers but influences the behavior of the agent by limiting the set of acceptable prizes in her action scheme. This is a crucial difference to our model, in which $\mathcal{P}$ cannot forbid certain prizes, but rather set up payments for each possible prize beforehand. A version of the delegated Pandora’s Box problem with transfers is studied in [5]. Here, the total exploration costs can either be shared between $\mathcal{P}$ and $\mathcal{A}$ by a fixed ratio, or the principal chooses the ratio as part of her strategy. This approach is substantially different from our contract setting with hidden actions. In the delegated variant, $\mathcal{P}$ can lower the opening costs of boxes for $\mathcal{A}$ to a value $c_{i}^{\prime}\leq c_{i}$ and pay the difference by herself, which requires direct observation of $\mathcal{A}$’s action (opening behavior). In contrast, in the contract setting, $\mathcal{A}$’s action is hidden from $\mathcal{P}$, and payments from $\mathcal{P}$ to $\mathcal{A}$ can only depend on the outcome selected in the end, not on the action of $\mathcal{A}$. Moreover, several key properties of delegation models are also adopted in [5]$\mathcal{P}$ commits to constraints on the set of acceptable outcomes, which is impossible in the contract setting. Finally, while in [4] the authors bound the delegation gap against the optimal non-delegated behavior of $\mathcal{P}$, we study the design of optimal contracts.
@@ 2 Preliminaries
$\mathcal{P}$ would like to motivate $\mathcal{A}$ to open and inspect boxes in order to find and select a good prize (for $\mathcal{P}$). $\mathcal{P}$ cannot observe the actions taken by $\mathcal{A}$ (i.e., which boxes were opened in which order), only the final selected box with the prize inside is revealed. Instead, $\mathcal{P}$ can commit to an exploration contract (or simply, contract) $T$: For each selected prize $(i,j)$, she specifies an amount $t_{ij}$ that she pays to $\mathcal{A}$. Then the final utility of $\mathcal{P}$ is $b_{ij}-t_{ij}$.
The task of the agent $\mathcal{A}$ is to open and inspect boxes to find and select at most one prize from an opened box. Formally, a policy specifies in every situation, based on the previously seen prizes, which box to open next, if any, or otherwise which prize to select from an opened box. Given the contract $T$, she is facing the classic Pandora’s Box problem [36]: She obtains a utility of $a_{ij}+t_{ij}$ for the selected prize (00 if no prize is selected) minus the opening cost $c_{i}$ for all opened boxes. The policies for $\mathcal{A}$ that maximize the expected total utility, called $\mathcal{A}$-optimal policies in the following, are well-known to be precisely the following index policies [36]: For each box $i\in[n]$ compute an index or fair cap $\smash{\varphi_{i}}$ such that
@@
The boxes are considered in non-increasing order of fair caps. Suppose box $i$ is the next unopened box in this order, and the best prize $\mathcal{A}$ has found thus far is $v$. If no box has been opened, then $v=0$, and otherwise $v=a_{i^{\prime}j^{\prime}}+t_{i^{\prime}j^{\prime}}$ for some $(i^{\prime},j^{\prime})$ where $i^{\prime}$ is a box that has already been opened, and prize $j^{\prime}$ has been found in it. There are three cases:
  • If $v<\smash{\varphi_{i}}$, $\mathcal{A}$ opens box $i$,
  • if $v=\smash{\varphi_{i}}$, $\mathcal{A}$ is indifferent between opening box $i$ and stopping opening boxes,
  • if $v>\smash{\varphi_{i}}$, $\mathcal{A}$ stops opening boxes.
In particular, any box $i$ with $\smash{\varphi_{i}}<0$ is never considered for opening. For each $(i,j)\in[n]\times[m]$ we define the capped value $\kappa_{i}=\min\{a_{ij}+t_{ij},\smash{\varphi_{i}}\}$. A policy as described above always selects the prize in box $i^{*}\in\arg\max_{i}\kappa_{i}$, i.e., the prize with the largest capped agent value.
Note that there may be several policies that are $\mathcal{A}$-optimal. Specifically,
@@
Lemma 1.
Given an exploration contract $T$, an optimal policy can be computed in polynomial time.
We study contracts for $\mathcal{P}$ that steer the index policy executed by $\mathcal{A}$ towards good outcomes for $\mathcal{P}$. We explore both linear and general contracts. A linear contract is given by a single number $\alpha\in[0,1]$, and $t_{ij}=\alpha\cdot b_{ij}$. Linear contracts are popular because of their simplicity, but they often suffer from substantial limitations of the achievable revenue. As such, we also explore general contracts, in which we only require every payment to be non-negative and bounded by the value obtained by $\mathcal{P}$, i.e., $0\leq t_{ij}\leq b_{ij}$ for every $i\in[n]$ and $j\in[m]$. A (linear) contract $T$ that, among all possible (linear) contracts, achieves maximum expected utility for $\mathcal{P}$ when $\mathcal{A}$ executes an optimal policy under $T$, is called optimal (linear) contract.
@@ 3 Optimal Linear Contracts
Observe that the critical values indeed have the property that, for any two such values $\alpha^{c}_{1},\alpha^{c}_{2}$, the set of $\mathcal{A}$-optimal policies is constant within $(\alpha^{c}_{1},\alpha^{c}_{2})$. Note that any $\mathcal{A}$-optimal policy is also $\mathcal{A}$-optimal under both $\alpha^{c}_{1}$ and $\alpha^{c}_{2}$, but there are potentially additional $\mathcal{A}$-optimal policies under $\alpha^{c}_{1}$ or under $\alpha^{c}_{2}$. Together with Observation 1, this fact implies that the optimal contract is a critical value.
The following auxiliary lemma helps us with analyzing the number of critical values.
@@
By Lemma 1, we can compute an optimal contract for any given critical value in polynomial time. The expected utility of $\mathcal{P}$ under this contract can also be computed in polynomial time. Thus, we can enumerate all critical values and take the best contract with respect to the expected utility for $\mathcal{P}$, implying Theorem 1.
@@ 4 Optimal General Contracts
In this section, we design optimal general (i.e., not necessarily linear) contracts in several natural settings.
4.1 No Intrinstic Agent Value
@@
As we observed above, an optimal contract cannot yield more utility for $\mathcal{P}$ than an optimal policy $\pi^{*}$ that she can apply by herself (paying the costs herself). Therefore, the following lemma implies Theorem 2.
@@
Recall that the following is an optimal policy $\pi^{*}$ for $\mathcal{P}$ (when she pays the costs herself). Define fair caps $\smash{\varphi^{\mathcal{P}}_{i}}$ for each box $i\in[n]$ such that
$\sum_{j\in[m]}p_{ij}\max\{0,b_{ij}-\smash{\varphi^{\mathcal{P}}_{i}}\}=c_{i}.$
Then $\mathcal{P}$ opens boxes in non-increasing order of their fair caps as long as the best prize observed so far does not exceed the fair cap of the next box.
@@
Consequently, $\mathcal{A}$ is faced with an instance of Pandora’s Box problem where her expected prize in each box $i$ equals the opening cost. Thus, if $\mathcal{A}$ applies the ($\mathcal{A}$-optimal) index policy for the emerging instance, the fair caps for $\mathcal{A}$ are given by $\smash{\varphi^{\mathcal{A}}_{i}}=0$ for all $i\in[n]$.

### v2

@@ 1 Introduction
In many real-world situations, e.g., the search for a job candidate or house, a decision maker is faced with the choice between different alternatives of a-priori unknown value, which can be explored at some cost. A large class of such settings can be modeled by the famous Pandora’s box model proposed and analyzed by Weitzman [39]. In this model, there are $n$ alternatives or boxes. The value in each box is distributed independently according to a known probability distribution. There is a known additive function for the exploration cost of the boxes. The goal is to maximize the value of the selected box minus the cost of explored boxes. Starting from [33], this problem has recently received a lot of attention at the intersection of economics and computer science (see, e.g., [38, 26, 8, 10, 6, 25] and a recent survey [7]).
@@
Due to missing qualifications or time constraints, in many markets the aforementioned exploration tasks are not executed directly by the decision maker. Rather, a decision maker (called principal $\mathcal{P}$ in the following) can delegate the exploration to an agent $\mathcal{A}$, whose incentives are potentially misaligned with that of $\mathcal{P}$.
Indeed, suppose the principal intends to buy a house. They might know an inspection cost and have some prior knowledge about the value of each house. However, they might not be qualified to determine the exact market value by inspecting it. They decide to delegate the search to a real-estate agent. The agent can have an individual valuation for each house (e.g., a provision that is paid internally for selling the house). This individual value might also depend on the condition of the house, which is revealed only after inspection. Similar situations arise, e.g., when the principal wants to invest in a financial product and delegates this search to a financial agent.
In these settings, as a consequence of delegating the search, $\mathcal{P}$ can merely observe the outcome of the search, not the other actions taken by $\mathcal{A}$. This leads to a possibly undesirable outcome for $\mathcal{P}$, even when they have the power to commit to accepting or rejecting certain outcomes [5].
@@
A natural approach to align the incentives of an agent with the goals of the principal in such hidden-action settings are contracts. A contract allows for utility transfers to be made from the principal to the agent as a function of the outcome of the delegated task. While contract theory is a well-established [9] and celebrated [35] area in economics, the interest in algorithmic contract design has surged only relatively recently, see e.g., [20, 21, 17, 14, 27].
@@
In this work, we initiate the study of algorithmic contract design for the Pandora’s Box problem. Specifically, we consider the following problem (for a fully formal definition, see Section 2): There are $n$ boxes, the $i$-th of which contains a prize, drawn independently from a known probability distribution, and can be opened at known cost $c_{i}$. The $j$-th prize from box $i$ has some utilities $a_{ij}$ and $b_{ij}$ for $\mathcal{A}$ and $\mathcal{P}$, respectively. Based on that information, $\mathcal{P}$ commits to a contract, specifying a transfer $t_{ij}\geq 0$ for each prize, where we assume non-negative transfers (limited liability). Given such a contract, $\mathcal{A}$ faces an instance of the classic Pandora’s Box Problem, which they solve optimally. Assuming $\mathcal{A}$ selects prize $j$ from box $i$, the resulting final utility of $\mathcal{P}$ is $b_{ij}-t_{ij}$. The goal is to efficiently compute a contract that maximizes the expectation of the latter quantity. Notably, ties in the optimization of $\mathcal{A}$ are broken in favor of $\mathcal{P}$, which is a standard assumption that makes the maximization problem well-defined.
@@
With a solution of $\mathcal{A}$’s problem at hand, the main trade-off inherent to an optimal contract can be understood as follows: On the one hand, transfers manipulate the fair caps, the resulting exploration order, and the acceptance decisions of $\mathcal{A}$ in such a way that leads to a more favorable outcome for $\mathcal{P}$. On the other hand, transfers make $\mathcal{P}$ lose utility since they are costly.
We start by considering linear contracts [30, 20], a class of contracts that has been particularly successful due to its simplicity. Here, a contract is fully described by a number $\alpha\in[0,1]$, which defines transfers $t_{ij}=\alpha\cdot b_{ij}$. We identify a polynomial-time computable set of critical values of $\alpha$ at which the set of $\mathcal{A}$-optimal policies (i.e., the set of policies that satisfy the above description) changes. By reasoning about the behavior of $\mathcal{P}$’s utility between any two such values, we can argue that the optimal contract needs to be a critical value. This type of argument is reminiscent of arguments from the recent literature (e.g., [17, 22, 16]), but here it requires special care. In particular, due to the tie breaking in favor of $\mathcal{P}$, which—as discussed above—is very relevant here, we need to keep track of the set of $\mathcal{A}$-optimal policies.
@@
The final case we consider has a more general value structure, but we compromise on the boxes having different distributions. Specifically, we consider instances with only single positive prize for $\mathcal{P}$, and all distributions are identical. Still, transfers can be made depending on the box from which the prize originates. Again, the complexity is reduced by the fact that only a single payment per box has to be decided. Through an intricate sequence of exchange arguments and the help of continuization, we establish that there is an optimal contract with a simple structure: (i) In a first phase, the prize with positive principal value gets immediately accepted, and all transfers are identical; (ii) all transfers for boxes in the second phase are also identical. These conditions allow to find the optimal contract by enumerating a polynomial-sized set of contracts.
@@
Overview.
After a review of related work in the subsequent section, we formally introduce our problem in Section 2 along with preliminary observations. Section 3 treats linear contracts. Section 4 deals with general contracts in several special cases (no agent value in Section 4.1 and binary boxes in Section 4.2). For spatial reasons, general contracts for i.i.d. boxes with single positive prize are discussed in Appendix C. Missing proofs are provided in Appendices A and B.
@@ 1.1 Related Work
Contract design has recently gained a lot of attention from a computational point of view. Duetting et al. [20] advocated the study of linear contracts as an alternative to the more complicated (and sometimes unintuitive) general contracts. They show robustness results and give parameterized approximation guarantees w.r.t. optimal (general) contracts. For settings in which the outcome space is succinctly represented, hardness results are shown in [21]. In a similar but different approach, combinatorial contracts [17] let the agent choose a subset of actions which stochastically determines the (binary) outcome. Important recent work in the area of combinatorial contracts includes [22], [16]. Contracts were also studied with multiple agents [18, 16], ambiguity [19], and private types [2] with connections to classic mechanism design [1]. More generally, contracts have been of interest in specific application domains, such as classification in machine learning [36].
Most closely related is prior work by Postl [34], who analyzes our contract-design problem for two boxes without intrinsic agent value. The paper provides an explicit characterization of the chosen box in the optimal contract, but neither considers algorithmic aspects nor the reduction to the standard Pandora’s Box problem we discover in Sec. 4.1. Concurrent to our work, Ezra et al [23] study the special case without agent value. In this special case, they also give a polynomial-time algorithm for linear contracts (without discussing the issues of tie-breaking in favor of $\mathcal{P}$). For general contracts, the paper studies a technical approach when there is a constant number of prizes. Finally, the authors show NP-hardness for correlated distributions in the boxes.
@@ 2 Preliminaries
$\mathcal{P}$ would like to motivate $\mathcal{A}$ to open and inspect boxes in order to find and select a good prize (for $\mathcal{P}$). $\mathcal{P}$ cannot observe the actions taken by $\mathcal{A}$ (i.e., which boxes were opened in which order), only the final selected box with the prize inside is revealed. Instead, $\mathcal{P}$ can commit to an exploration contract (or simply, contract) $T$: For each selected prize $(i,j)$, she specifies an amount $t_{ij}\in[0,b_{ij}]$ [^1] We make the natural assumption that $t_{ij}\leq b_{ij}$, i.e., every transfer in the contract is bounded by the prize value of $\mathcal{P}$. For all scenarios we study here (linear contracts, no agent value, binary boxes, i.i.d. with single positive prize for $\mathcal{P}$) it is straightforward to see that this assumption is w.l.o.g. – there is an optimal contract that satisfies $t_{ij}\leq b_{ij}$ for all $(i,j)\in[n]\times[m]$. It is a very interesting open problem to prove that this holds beyond the cases we study here. that she pays to $\mathcal{A}$. Then the final utility of $\mathcal{P}$ is $b_{ij}-t_{ij}$.
Our model is an instance of the standard principal-agent model, where the hidden action of the agent $\mathcal{A}$ is the policy to inspect and select from the boxes. Specifically, the task of $\mathcal{A}$ is to open and inspect boxes to find and select at most one prize from an opened box. Formally, a policy specifies in every situation, based on the previously seen prizes, which box to open next, if any, or otherwise which prize to select from an opened box. Given a contract $T$, she is facing the classic Pandora’s Box problem [39]: She obtains a utility of $a_{ij}+t_{ij}$ for the selected prize (00 if no prize is selected) minus the opening cost $c_{i}$ for all opened boxes. Weitzman [39] shows that every optimal policy of $\mathcal{A}$ that maximizes his expected total utility (called $\mathcal{A}$-optimal policy in the following) is in the form of the following index policy: For each box $i\in[n]$ compute an index or fair cap $\smash{\varphi_{i}}$ such that
@@
The boxes are considered in non-increasing order of fair caps. Suppose box $i$ is the next unopened box in this order, and the best prize $\mathcal{A}$ has found thus far is $v$. If no box has been opened, then $v=0$, and otherwise $v=a_{i^{\prime}j^{\prime}}+t_{i^{\prime}j^{\prime}}$ for some $(i^{\prime},j^{\prime})$ where $i^{\prime}$ is a box that has already been opened, and prize $j^{\prime}$ has been found in it. There are three cases: (1) $v<\smash{\varphi_{i}}$: $\mathcal{A}$ opens box $i$; (2) $v=\smash{\varphi_{i}}$: $\mathcal{A}$ is indifferent between opening box $i$ and stopping opening boxes; and (3) $v>\smash{\varphi_{i}}$: $\mathcal{A}$ stops opening boxes. Note that any box $i$ with $\smash{\varphi_{i}}<0$ is never opened. For each $(i,j)\in[n]\times[m]$ we define the capped value $\kappa_{i}=\min\{a_{ij}+t_{ij},\smash{\varphi_{i}}\}$. A policy as described above always selects the prize in box $i^{*}\in\arg\max_{i}\kappa_{i}$, i.e., the prize with the largest capped agent value.
Even under these conditions, there may still be several $\mathcal{A}$-optimal policies. Specifically,
@@
Theorem 1.
Given a contract $T$, an optimal policy can be computed in polynomial time.
We study contracts for $\mathcal{P}$ that steer the index policy executed by $\mathcal{A}$ towards good outcomes for $\mathcal{P}$. We explore both linear and general contracts. A linear contract is given by a single number $\alpha\in[0,1]$, and $t_{ij}=\alpha\cdot b_{ij}$. Linear contracts are popular because of their simplicity, but they often suffer from substantial limitations of the achievable revenue. As such, we also explore general contracts, in which we only require every payment to be non-negative $t_{ij}\geq 0$ for every $i\in[n]$ and $j\in[m]$. A (linear) contract $T$ that, among all possible (linear) contracts, achieves maximum expected utility for $\mathcal{P}$ when $\mathcal{A}$ executes an optimal policy under $T$, is called optimal (linear) contract.
@@ 3 Optimal Linear Contracts
Observe that the critical values indeed have the property that, for any two such values $\alpha^{c}_{1},\alpha^{c}_{2}$, the set of $\mathcal{A}$-optimal policies is constant within $(\alpha^{c}_{1},\alpha^{c}_{2})$. Note that any $\mathcal{A}$-optimal policy is also $\mathcal{A}$-optimal under both $\alpha^{c}_{1}$ and $\alpha^{c}_{2}$, but there are potentially additional $\mathcal{A}$-optimal policies under $\alpha^{c}_{1}$ or under $\alpha^{c}_{2}$. Together with Observation 1, this implies that the optimal contract is a critical value.
The following auxiliary lemma supports the analysis of the number of critical values.
@@
By Theorem 1, we can compute an optimal contract for any given critical value in polynomial time. The expected utility of $\mathcal{P}$ under this contract can also be computed in polynomial time. Thus, we can enumerate all critical values and take the best contract with respect to the expected utility for $\mathcal{P}$, implying Theorem 2.
@@ 4 Optimal General Contracts
4.1 No Intrinsic Agent Value
@@
As observed above, an optimal contract cannot yield more utility for $\mathcal{P}$ than an optimal policy $\pi^{*}$ that she can apply by herself (paying the costs herself). Hence, the next lemma implies Theorem 3.
@@
Recall that the following is an optimal policy $\pi^{*}$ for $\mathcal{P}$ (when she pays the costs herself). Define fair caps $\smash{\varphi^{\mathcal{P}}_{i}}$ for each box $i\in[n]$ such that $\sum_{j\in[m]}p_{ij}\max\{0,b_{ij}-\smash{\varphi^{\mathcal{P}}_{i}}\}=c_{i}$. Then $\mathcal{P}$ opens boxes in non-increasing order of their fair caps as long as the best prize observed so far does not exceed the fair cap of the next box.
@@
Consequently, $\mathcal{A}$ is faced with an instance of the Pandora’s Box problem where her expected prize in each box $i$ equals the opening cost. Thus, if $\mathcal{A}$ applies the ($\mathcal{A}$-optimal) index policy for the emerging instance, the fair caps for $\mathcal{A}$ are given by $\smash{\varphi^{\mathcal{A}}_{i}}=0$ for all $i\in[n]$.